OpenStudy (anonymous):

A line passes through the point (–7, 5) and has a slope of 1/2. Which is another point that the line passes through? A. (–13, 9) B. (–9, 13) C. (9, 13) D. (13, 9)

OpenStudy (anonymous):

Please Someone Help Me!

OpenStudy (imstuck):

I'm doing it now...brb

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

It should be B.

OpenStudy (anonymous):

i put B first and it was wrong.

OpenStudy (anonymous):

Oh I'm so sorry I messed up

OpenStudy (anonymous):

Slope equation is y2-y1/x2-x1= slope You are given two points already so you can plug and chug.

OpenStudy (anonymous):

Oh Okay Thank You

OpenStudy (anonymous):

It's C actually try that, (5-13)/(-7-9)= 1/2 -8/-16= 1/2 Simplify and it should be 1/2

OpenStudy (anonymous):

Thank You